# Localized states, topology and anomalous Hall conductivity on a 30 degrees twisted bilayer honeycomb lattice

### v3

@@ II.2 Multifractal properties
To explore multifractal properties of our model, we plot average fractal dimension and its standard deviation for all states against the system size and a few values of $q$ (Fig. 9). At zero $t_{inter}$, we observe that as the system size increases, the average fractal dimension of all states in the system approaches $1$ and its standard deviation approaches 0. In other words, we confirm that in both topological and non-topological weakly coupled phases, the bulk eigenstates are fully extended as they are expected to be in a fully periodic system.
@@
Figure 9: Average fractal dimensions $D_{ave}$ and their standard deviations $\Delta D$ over all states in a system plotted against its size $L$. We choose a list of $q=0.5,0.75,1.25,1.5,2.0,3.0,4.0$. In the weakly coupled phase, $D_{ave}$ converges to $1$, and $\Delta D$ converges to 0 - this means that all states are extended. In the strongly coupled phase, both $D_{ave}$ and $\Delta D$ converge to values between 0 and 11 - the system is multifractal.
@@
Figure 10: Fractal dimensions $D$ of the states with the largest energies plotted against $1/\log(L_{x})$ for $q=0.5,0.75,1.25,1.5,2.0,3.0,4.0$. The wavefunctions were obtained using Lanczos diagonaliztion. We see that for $t_{inter}=0$, $D$ approaches $1$ with increasing $L_{x}$, but does not approach 11 for $t_{inter}\neq 0$.
@@
Characterizing multifractal properties at non-zero $t_{inter}$ turns out to be challenging because finite size corrections to fractal dimensions scale as $\log(L) 11$1One can see this just by writing a wavefunctions as $\psi\sim C/\sqrt{L}$ and plugging into the Eq. 5, and because of that it is impossible to obtain reliable results for any realistic system size, which can be handled numerically. For this reason, we cannot say confidently whether average fractal dimensions approach $1$ at $t_{inter}=1$. However, when we plot average fractal dimensions and their deviations as a function of the system size in the ’strongly coupled’ phase ($t_{inter}=10$), we observe that the former converge to values smaller than $1$, and the latter converge to values greater than 0. Furthermore, we can see that as the system size grows, average fractal dimensions converge to different values for various $q$. Thus we can say fairly confidently that eigenstates in the strongly coupled phase are multifractal - they are neither extended, nor localized.
@@
To complement our understanding of multifractaility in our system, we also study it using Lanczos diagonalization. This method makes it possible to consider much larger system sizes (up to $L_{x}\approx 800$). Specifically, for each $L_{x}$, we obtain eigenstates with the largest energies and compute their fractal dimensions (see Fig. 10). As previously, we can see that for $t_{inter}=0$, i.e. for periodic system, fractal dimension of the ‘top’ state approaches $1$ with increasing system size. On the other hand, for $t_{inter}\neq 0$, we can see that the fractal dimension of the ‘top’ state does not approach $1$. However, despite reaching huge lattice sizes, we are still not able to ensure that they are sufficient to make predictions about infinite-size properties. For instance, at $m=1.5$, $t_{inter}=10$, we observe that the wavefunction of the ‘top’ state remains rotationally-symmetric for sizes up to $L_{x}\approx 420$, whereas at larger sizes, starts getting localized near one of the corners (see Fig. 11). We believe that this is yet another manifestation of the fact that in a quasicrystalline system, continuum limit is reached extremely slowly, and even system with the largest sizes that can be handled by modern computers may still be far from it. In the future, it would be of interest to find means to study multifractality in our model analytically.
@@
Figure 11: Wavefunctions of the states with highest energies obtained by Lanczos diagonalization for $m=1.5$, $t_{inter}=10$, $Lx=350$ (11(a)) and $L_{x}=450$ (11(b)).
@@ II.3.1 Topological entanglement entropy
To study entanglement properties of a many body system, one has to partition it into two subsystems in coordinate space. For a subsystem with density matrix $\rho_{A}$, the entanglement entropy is defined as $S=-\mathrm{tr}\rho_{A}\log\rho_{A}$. If a 2D system has smooth boundary, its entanglement entropy scales linearly with system size $L$ as
@@
To extract the topological contribution $\gamma$, we compute the difference between entanglement entropies of several systems in such a way that their leading term, which is linear over $L$, as well as subleading contributions due to corners exactly cancel out. To do this, we choose a configuration of four systems shown on the Fig. 12. We compute entanglement entropies of each subsystem using the method suggested in Refs. [20, 21, 23], which we briefly describe in Sec. A.
Figure 12: A scheme of partitioning the system, which we use to compute the topological contribution $\gamma$ to the entanglement entropy. Each of the circular curves has a radius of $L_{x}/2$. We compute entanglement entropy for each of the figures and then use this substraction scheme to isolate the contribution from the topological term $\gamma$.
@@
We also compute $\gamma$ for large $t_{inter}$ at $m=0$, assuming that the Fermi level $E_{F}$ lies within the energy gap, which we previously found on the Figs 4(a), 4(b), 4(c). We obtain that at large system size, $\gamma$ becomes close to zero (see Fig. 14). This suggests that our ‘strongly coupled’ gapped system is not topological.
@@
Figure 13: Topological entanglement entropy $\gamma$ plotted against $t_{inter}$ (13(a)) for $m=0,2$ and against $m$ (13(b)) for $t_{inter}=0.0,1.0,10.0$. Here we assume $E_{F}=0$. It is equal to a non-zero constant in the topological gapped phase and zero in the non-topological phase
Figure 14: Topological entanglement entropy $\gamma$ plotted against $t_{inter}$ for $m=0$, in the case when the Fermi level lies within the bulk energy gap of the ‘strongly coupled’ system. We can see that with increasing system size $L_{x}$, it approaches zero.
@@ II.3.2 Local Chern marker
It was argued in Ref [11], that Chern number is not well-defined for a system of finite size because if one defines it in the same way as the corresponding infinite-size expression, the answer becomes exactly zero. Instead it was suggested that topological properites of a system can be characterized empirically using ’local Chern marker’ defined in the following way. First, one can define a projector to the subspace of filled states as $P(r_{i},r_{j})=\sum\limits_{E_{\lambda}<E_{F}}\psi_{\lambda}(r_{i})\psi^{\dagger}_{\lambda}(r_{j})$. Next, one may project the coordinate operators as $\tilde{X}(r_{i},r_{j})=\sum\limits_{r_{k}}P(r_{i},r_{k})x_{k}P(r_{k},r_{j})$ and $\tilde{Y}(r_{j},r_{i})=\sum\limits_{r_{k^{\prime}}}P(r_{j},r_{k^{\prime}})y_{k^{\prime}}P(r_{k^{\prime}},r_{i})$. The local Chern marker is defined as
$C(r_{i})=\frac{2\pi i}{S}\sum_{r_{j}}\left[\tilde{X}(r_{i},r_{j}),\tilde{Y}(r_{j},r_{i})\right],$ (7)
@@
where $S$ is an average area per each lattice site on the lattice. One may check that for conventional Haldane model on a finite lattice, this local Chern marker inside the bulk is equal to the value of Chern invariant in the corresponding continuum model, whereas near the edges, its value is very large and has opposite magnitude in such a way that its sum over all lattice sites is zero: $\sum\limits_{i}C(r_{i})=0$.
@@
We plot the local Chern marker of the $30^{\circ}$ twisted bilayer Haldane model for various values of $t_{inter}$ on Fig. 15(a). One may see that in the ’weakly coupled’ phases, it behaves in the same way as in the Haldane model on a monolayer. Specifically, in the ’weakly coupled topological phase’, its bulk value remains the same and equal to the value in the case of zero interlayer coupling. Similarly, in the ’weakly coupled non-topological’ phase, its bulk value is equal to zero. On the other hand, at larger values of $t_{inter}$, the local Chern marker starts behaving differently from the case of periodic lattice. Specifically, the local Chern marker no longer has a form of constant value in the bulk and large, but opposite value at the edge. Instead, regions with $C(r_{i})$ values of opposite signs start emerging within the bulk of the system, and sufficiently far away from the transition point, these regions become indistinguishable from the edges of the system. In other words, at large $t_{inter}$, the local Chern marker $C(r_{i})$ strongly fluctuates within the bulk. so that there is no distinction between the bulk and the edges as in a crystalline topological insulator.
Finally, we point out that the behavior of the local Chern marker demonstrates a ‘hidden’ phase transition at a value of $\tilde{t}_{inter}$ smaller than the bulk gap closing (see Fig. 15(a)). Specifically, at $t_{inter}<\tilde{t}_{inter}$, the local Chern marker value inside the bulk is constant across the lattice and equal to the value corresponding to the limit of zero interactions. However, at $t_{inter}>\tilde{t}_{inter}$, the local Chern marker starts fluctuating within the lattice, and its average value starts changing. We emphasize that such a phase transition cannot be observed by looking at the energy spectrum alone. Thus we conclude that the local Chern marker can be used to detect novel phase transitions within quasicrystals.
@@
Figure 15: 15(a) Local Chern marker plotted for lattices with L<x,y<L-L<x,y<L at various values of $m$, $t_{inter}$. One can see that in the bulk of a topological phase, it is approximately equal to Chern invariant, and in the non-topological phase, it is equal to zero. In the multifractal phase, it fluctuates within the lattice. 15(b) Averaged local Chern marker $C_{ave}$ and its standard deviation $\Delta C$ taken over the inner part of the lattice with L/2<x,y<L/2-L/2<x,y<L/2. It remains constant and approximately equal to the bulk Chern invariant in the weakly coupled phase, but starts fluctuating at large $t_{inter}$.
@@ II.3.3 Anomalous Hall conductivity
We start our analysis from introducing electric current of our model. Specifically, if a tight-binding Hamiltonian for a given system has hoppings $H_{ij}$ between sites $i,j$, then in the presence of electric field they change according to Peierls substitution and thus become $H_{ij}e^{\frac{ie}{\hbar}\int\limits_{r_{i}}^{r_{j}}\vec{A}\vec{dr}}$, where $\vec{A}$ is a vector potential. If we expand these hoppings to the first order in $\vec{A}$, we obtain that the electric current operator has the following expression
@@
Further, one can compute a response to applied electric field by using an expression for the electric current (8) and applying Kubo formula. Anomalous Hall conductivity would be its antisymmetric part, and thus it would be given by the Eq. (16). After lengthy, but straightforward calculations (specifically writing electrons Green functions and performing Matsubaba summation), the expression for anomalous Hall conductivity can be brought to the following form
$\sigma^{(xy)}_{ij}$ $=$ $-\frac{e^{2}}{2\hbar^{2}}\sum\limits_{f,e}\sum\limits_{k,l}\frac{1}{(E_{f}-E_{e})^{2}}$
@@
In this equation, $i,j$ are lattice sites between which the electric current is computed, $k,l$ are other pairs of sites over which summation is performed, $E,\psi$ are eigenenergies and eigenfunctions of the model, indices $f,e$ numerate filled/empty states respectively.
Let us emphasize once again, that in the Eq. (II.3.3), $\sigma_{ij}$ describes electric current through a bond between the lattice sites $i,j$ (we assumed that the electric field is uniform). However, one can sum electric currents through all bonds next to a given lattice site $i$ and consider $\sigma_{i}=\sum\limits_{j}\sigma_{ij}$. For a crystalline system, one can take a step further and split the lattice site index $i$ into two indices $i^{\prime},\alpha$ numerating unit cells and sublattice degrees of freedom respectively. The well-known integer value of the anomalous Hall conductivity is obtained after an additional summation over $\alpha$. A simple explanation of the latter is as follows: for an infinite crystalline system one may transform any dependency over unit cell index $i^{\prime}$ into momentum representation, and the well-known integer AHE occurs at zero momentum. However, its expression in terms of Berry curvature still contains summation over sublattice degrees of freedom. Nevertheless, on a honeycomb lattice, the two sublattice degrees of freedom are equivalent to each other, and therefore the AHE for each of them would be the same, i.e. half-integer.
@@
Figure 16: 16(a) Hall conductivity $\sigma_{i}$ (in units of $e^{2}/\hbar^{2}$) at each site $i$ plotted for lattices with L<x,y<L-L<x,y<L at various values of $m$, $t_{inter}$. One can see that in the bulk of a topological phase, it is approximately equal to $0.5$ - Chern invariant divided by the number of sites in a unit cell in each layer, and in the non-topological phase, it is equal to zero. In the multifractal phase, it fluctuates within the lattice. 16(b) Averaged $\sigma_{i}$ summed over the four sublattice degrees of freedom ($A/B$ and top/bottom) and its standard deviation taken over the inner part of the lattice with L/2<x,y<L/2-L/2<x,y<L/2. It remains constant and approximately equal to the bulk Chern invariant in the weakly coupled phase, but starts fluctuating at large $t_{inter}$.

### v4

@@ II.2 Multifractal properties
To explore multifractal properties of our model, we plot average fractal dimension and its standard deviation for all states against the system size (once again we assume that the system forms a square with $L_{x}=L_{y}$) and a few values of $q$ (Fig. 10). At zero $t_{inter}$, we observe that as the system size increases, the average fractal dimension of all states in the system approaches $1$ and its standard deviation approaches 0. In other words, we confirm that in both topological and non-topological weakly coupled phases, the bulk eigenstates are fully extended as they are expected to be in a fully periodic system.
@@
Figure 10: Average fractal dimensions $D_{q,ave}$ and their standard deviations $\Delta D_{q}$ over all states in a system plotted against its size $L_{x}$ (we assume $L_{x}=L_{y}$). We choose a list of $q=0.5,0.75,1.25,1.5,2.0,3.0,4.0$. In the weakly coupled phase, $D_{q,ave}$ converges to $1$, and $\Delta D_{q}$ converges to 0 - this means that all states are extended. In the strongly coupled phase, both $D_{q,ave}$ and $\Delta D_{q}$ converge to values between 0 and 11 - the system is multifractal.
@@
Figure 11: Fractal dimensions $D_{q}$ of the states with the largest energies plotted against $1/\log(L_{x})$ for $q=0.5,0.75,1.25,1.5,2.0,3.0,4.0$. We assume $L_{x}=L_{y}$. The wavefunctions were obtained using Lanczos diagonaliztion. We see that for $t_{inter}=0$, $D_{q}$ approaches $1$ with increasing $L_{x}$, but does not approach 11 for $t_{inter}\neq 0$.
@@
Characterizing multifractal properties at non-zero $t_{inter}$ turns out to be challenging because finite size corrections to fractal dimensions scale as $1/\log(L)$ for a system with a linear size $L$  [^1] One can see this just by writing a wavefunctions as $\psi\sim C/\sqrt{L}$ and plugging into the Eq. 5, and because of that it is impossible to obtain reliable results for any realistic system size, which can be handled numerically. For this reason, we cannot say confidently whether average fractal dimensions approach $1$ at $t_{inter}=1$. However, when we plot average fractal dimensions and their deviations as a function of the system size in the ’strongly coupled’ phase ($t_{inter}=10$), we observe that the former converge to values smaller than $1$, and the latter converge to values greater than 0. Furthermore, we can see that as the system size grows, average fractal dimensions converge to different values for various $q$. Thus we can say fairly confidently that eigenstates in the strongly coupled phase are multifractal - they are neither extended, nor localized.
@@
To complement our understanding of multifractaility in our system, we also study it using Lanczos diagonalization. This method makes it possible to consider much larger system sizes (up to $L_{x,y}\approx 800$). Specifically, for each $L_{x,y}$, we obtain eigenstates with the largest energies and compute their fractal dimensions (see Fig. 11). As previously, we can see that for $t_{inter}=0$, i.e. for periodic system, fractal dimension of the ‘top’ state approaches $1$ with increasing system size. On the other hand, for $t_{inter}\neq 0$, we can see that the fractal dimension of the ‘top’ state does not approach $1$. However, despite reaching huge lattice sizes, we are still not able to ensure that they are sufficient to make predictions about infinite-size properties. For instance, at $m=1.5$, $t_{inter}=10$, we observe that the wavefunction of the ‘top’ state remains rotationally-symmetric for sizes up to $L_{x,y}\approx 420$, whereas at larger sizes, starts getting localized near one of the corners (see Fig. 12). We believe that this is yet another manifestation of the fact that in a quasicrystalline system, continuum limit is reached extremely slowly, and even system with the largest sizes that can be handled by modern computers may still be far from it. In the future, it would be of interest to find means to study multifractality in our model analytically.
@@
Figure 12: $|\psi_{i}|^{2}$ of the states with highest energies obtained by Lanczos diagonalization for $m=1.5$, $t_{inter}=10$, $L_{x,y}=350$ (12(a)) and $L_{x,y}=450$ (12(b)).
@@ II.3.1 Topological entanglement entropy
To study entanglement properties of a many body system, one has to partition it into two subsystems in coordinate space. For a subsystem $A$ with density matrix $\rho_{A}$, the entanglement entropy is defined as $S=-\mathrm{tr}\rho_{A}\log\rho_{A}$. If a 2D system has smooth boundary, its entanglement entropy scales linearly with system size $L$ as
@@
To extract the topological contribution $\gamma$, we compute the difference between entanglement entropies of several subsystems in such a way that their leading term, which is linear over $L$, as well as subleading contributions due to corners exactly cancel out. To do this, we choose a configuration of four systems shown on the Fig. 13. We compute entanglement entropies of each subsystem using the method suggested in Refs. [9, 20, 2], which we briefly describe in Sec. B.
Figure 13: A scheme of partitioning a system, which we use to compute the topological contribution $\gamma$ to the entanglement entropy. Each of the circular curves has a radius of $L_{x}/2$. We compute entanglement entropy for each of the figures and then use this substraction scheme to isolate the contribution from the topological term $\gamma$.
@@
We also compute $\gamma$ for large $t_{inter}$ at $m=0$, assuming that the Fermi level $E_{F}$ lies within the energy gap, which we previously found on the Figs 5(a), 5(b), 5(c). We obtain that at large system size, $\gamma$ becomes close to zero (see Fig. 15). This suggests that our ‘strongly coupled’ gapped system is not topological.
@@
Figure 14: Topological entanglement entropy $\gamma$ plotted against $t_{inter}$ (14(a)) for $m=0,2$ and against $m$ (14(b)) for $t_{inter}=0.0,1.0,10.0$. Here we assume $E_{F}=0$. It is equal to a non-zero constant in the topological gapped phase and zero in the non-topological phase
Figure 15: Topological entanglement entropy $\gamma$ plotted against $t_{inter}$ for $m=0$, in the case when the Fermi level lies within the bulk energy gap of the ‘strongly coupled’ system. We can see that with increasing system size $L_{x}$ (we assume $L_{x}=L_{y}$), it approaches zero.
@@ II.3.2 Local Chern marker
It was argued in Ref [5], that Chern number is not well-defined for a system of finite size because if one defines it in the same way as the corresponding infinite-size expression, the answer becomes exactly zero. Instead it was suggested that topological properites of a system can be characterized empirically using ’local Chern marker’ defined in the following way. First, one can define a projector to the subspace of filled states as $P(\vec{r}_{i},\vec{r}_{j})=\sum\limits_{E_{f}<E_{F}}\psi_{f}(\vec{r}_{i})\psi^{\dagger}_{f}(\vec{r}_{j})$ (here the indices $i,j$ numerate the lattice sites, and the summation is performed over filled states numerated by an index $f$). Next, one may project the coordinate operators as $\tilde{X}(\vec{r}_{i},\vec{r}_{j})=\sum\limits_{\vec{r}_{k}}P(\vec{r}_{i},\vec{r}_{k})x_{k}P(\vec{r}_{k},\vec{r}_{j})$ and $\tilde{Y}(\vec{r}_{j},\vec{r}_{i})=\sum\limits_{\vec{r}_{k^{\prime}}}P(\vec{r}_{j},\vec{r}_{k^{\prime}})y_{k^{\prime}}P(\vec{r}_{k^{\prime}},\vec{r}_{i})$. The local Chern marker is defined as
$C(\vec{r}_{i})=\frac{2\pi i}{S}\sum_{\vec{r}_{j}}\left[\tilde{X}(\vec{r}_{i},\vec{r}_{j}),\tilde{Y}(\vec{r}_{j},\vec{r}_{i})\right],$ (7)
@@
where $S$ is an average area per each lattice site on the lattice. One may check that for conventional Haldane model on a finite lattice, this local Chern marker inside the bulk is equal to the value of Chern invariant in the corresponding continuum model, whereas near the edges, its value is very large and has opposite magnitude in such a way that its sum over all lattice sites is zero: $\sum\limits_{i}C(\vec{r}_{i})=0$.
@@
We plot the local Chern marker of the $30^{\circ}$ twisted bilayer Haldane model at $E_{F}=0$ for various values of $t_{inter}$ on Fig. 16(a). One may see that in the ’weakly coupled’ phases, it behaves in the same way as in the case of two non-coupled copies of Haldane model on a monolayer. Specifically, in the ’weakly coupled topological phase’, its bulk value remains the same and equal to the value in the case of zero interlayer coupling. Similarly, in the ’weakly coupled non-topological’ phase, its bulk value is equal to zero. On the other hand, at larger values of $t_{inter}$, the local Chern marker starts behaving differently from the case of periodic lattice. Specifically, the local Chern marker no longer has a form of constant value in the bulk and large, but opposite value at the edge. Instead, regions with $C(\vec{r}_{i})$ values of opposite signs start emerging within the bulk of the system, and sufficiently far away from the transition point, these regions become indistinguishable from the edges of the system. In other words, at large $t_{inter}$, the local Chern marker $C(\vec{r}_{i})$ strongly fluctuates within the bulk. so that there is no distinction between the bulk and the edges as in a crystalline topological insulator.
For completeness, we also plot the local Chern marker in the strongly coupled phase ($t_{inter}=10$) in the case when $E_{F}$ lies within the bulk gap (see Fig. 17). We can see once again that it strongly fluctuates within the bulk. Thus, we conclude that the strongly coupled gapped phase cannot be characterized by a well-defined behavior of the local Chern marker, which is in turn just another evidence that this phase is non-topological.
@@
Figure 16: 16(a) Local Chern marker plotted for lattices with $L_{x,y}<x,y<L_{x,y}$ at various values of $m$, $t_{inter}$ and $L_{x}=L_{y}=30$. One can see that in the bulk of a topological phase, it is approximately equal to Chern invariant, and in the non-topological phase, it is equal to zero. In the multifractal phase, it fluctuates within the lattice. 16(b) Averaged local Chern marker $C_{ave}$ and its standard deviation $\Delta C$ taken over the inner part of the lattice with $L_{x,y}/2<x,y<L_{x,y}/2$. It remains constant and approximately equal to the bulk Chern invariant in the weakly coupled phase, but starts fluctuating at large $t_{inter}$.
Figure 17: Local Chern marker $C(\vec{r}_{i})$ plotted for the system in the strongly coupled phase ($t_{inter}=10$) in the case when the Fermi level lies within the bulk energy gap. We assume $m=0.0$ and consider lattice sizes $L_{x,y}=20,30,40$. One can see that $C(\vec{r}_{i})$ fluctuates within the bulk, which in turn means that the system does not have well-defined topological properties.
@@ II.3.3 Anomalous Hall conductivity
We start our analysis from introducing electric current of our model. Specifically, if a tight-binding Hamiltonian for a given system has hoppings $H_{ij}$ between sites $i,j$, then in the presence of electric field they change according to Peierls substitution and thus become $H_{ij}e^{\frac{ie}{\hbar}\int\limits_{\vec{r}_{i}}^{\vec{r}_{j}}\vec{A}\vec{dr}}$, where $\vec{A}$ is a vector potential at the bond connecting the sites at positions $\vec{r}_{i}$ and $\vec{r}_{j}$. If we expand these hoppings to the first order in $\vec{A}$, we obtain that the electric current operator has the following expression
@@
Further, one can compute a response to applied electric field by using an expression for the electric current (8) and applying Kubo formula. Anomalous Hall conductivity would be its antisymmetric part, and thus it would be given by the Eq. (17). After lengthy, but straightforward calculations (specifically writing electrons Green functions and performing Matsubaba summation; see Sec. C for more details), the expression for anomalous Hall conductivity can be brought to the following form
$\sigma^{(xy)}_{ij}$ $=$ $-\frac{e^{2}}{2\hbar^{2}}\sum\limits_{f-filled\atop e-empty}\sum\limits_{k,l}\frac{1}{(E_{f}-E_{e})^{2}}$
@@
In this equation, the indices $i,j$ numerate the lattice sites between which the electric current is computed, and $k,l$ numerate the sites over which summation is performed. In addition, the summation is performed over filled states numerated by an index $f$ and empty states numerated by an index $e$. Specifically, $E_{f},E_{e}$ are energies of the filled and empty states respectively, and similarly $\psi_{f,i}\equiv\psi_{f}(\vec{r}_{i})$ and $\psi_{e,i}\equiv\psi_{e}(\vec{r}_{i})$ are the wavefunctions of the filled and empty states respectively at the lattice site $i$.
Let us emphasize once again, that in the Eq. (II.3.3), $\sigma_{ij}$ describes electric current through a bond between the lattice sites $i,j$ (we assumed that the electric field is uniform). However, one can sum electric currents through all bonds next to a given lattice site $i$ and consider $\sigma_{i}=\sum\limits_{j}\sigma_{ij}$. For a crystalline system, one can take a step further and split the lattice site index $i$ into two indices $i^{\prime},\alpha$ numerating unit cells and sublattice degrees of freedom respectively. The well-known integer value of the anomalous Hall conductivity is obtained after an additional summation over $\alpha$ (see Sec. D). A simple explanation of the latter is as follows: for an infinite crystalline system one may transform any dependency over unit cell index $i^{\prime}$ into momentum representation, and the well-known integer AHE occurs at zero momentum. However, its expression in terms of Berry curvature still contains summation over sublattice degrees of freedom. Nevertheless, on a honeycomb lattice, the two sublattice degrees of freedom are equivalent to each other, and therefore the AHE for each of them would be the same, i.e. half-integer.
@@
Figure 18: 18(a) Hall conductivity $\sigma_{i}$ (in units of $e^{2}/\hbar^{2}$) at each site $i$ plotted for lattices with $L_{x,y}<x,y<L_{x,y}$ at various values of $m$, $t_{inter}$ and $L_{x}=L_{y}=30$. One can see that in the bulk of a topological phase, it is approximately equal to $0.5$ - Chern invariant divided by the number of sites in a unit cell in each layer, and in the non-topological phase, it is equal to zero. In the multifractal phase, it fluctuates within the lattice. 18(b) Averaged $\sigma_{i}$ summed over the four sublattice degrees of freedom ($A/B$ and top/bottom) and its standard deviation taken over the inner part of the lattice with $L_{x,y}/2<x,y<L_{x,y}/2$. It remains constant and approximately equal to the bulk Chern invariant in the weakly coupled phase, but starts fluctuating at large $t_{inter}$.